# Resolving UnderEdit & OverEdit with Iterative & Neighbor-Assisted Model Editingthanks: Preprint under-review

### v2

@@ Abstract
Large Language Models (LLMs) are widely deployed in downstream tasks, but keeping their knowledge up-to-date via retraining or fine-tuning is often computationally expensive. Model editing provides a more efficient alternative by updating a targeted subset of parameters, which often follows the locate-and-edit paradigm. Despite this efficiency, existing methods are limited: edits may fail to inject knowledge (UnderEdit) or unintentionally disrupt unrelated neighboring knowledge (OverEdit). To address these challenges, we propose two complementary methods: iterative model editing, which applies successive edits to mitigate UnderEdit, and neighbor-assisted model editing, which incorporates neighboring knowledge during editing to reduce OverEdit. Our extensive experiments show that these techniques improve editing performance across multiple LLMs, algorithms, and benchmarks, reducing UnderEdit by up to 38 percentage points and OverEdit by up to 6, while remaining broadly applicable to any locate-and-edit method.
@@
Iterative & Neighbor-Assisted Model Editingthanks: Preprint under-review
Bhiman Kumar Baghel  Scott M. Jordan  Zheyuan Ryan Shi  Xiang Lorraine Li Department of Computer Science, University of Pittsburgh, PA, USA {bkb45, scott.jordan, ryanshi, xianglli}@pitt.edu
@@ 1 Introduction
LLMs have been widely used as repositories of factual and specialized knowledge (Petroni et al., 2020; Jiang et al., 2021; Roberts et al., 2020; Youssef et al., 2023). However, the world is constantly changing, with knowledge and information evolving rapidly, such as significant government policy changes and their wide impacts across various domains. Thus, it is essential for many NLP applications, such as text generation, question answering, and knowledge retrieval, to have models that can adapt to knowledge changes both effectively and efficiently. Re-training an LLM is resource-intensive (Patterson et al., 2021). Standard supervised fine-tuning is data hungry and less effective (Meng et al., 2023b). Model-editing, which directly modifies important model parameters for making the prediction, has emerged as a more efficient alternative for updating outdated information (Meng et al., 2023a, b; Li et al., 2024; Fang et al., 2025). These methods adopt a “locate-and-edit” approach, where they first identify the parameter locations associated with outdated knowledge and then update the parameters to enable the model to incorporate and predict the new knowledge.
@@
In summary, we propose solutions to two fundamental challenges in model editing: UnderEdit, where edits fail, and OverEdit, where neighboring knowledge is erroneously modified. We evaluate our approach using four “locate and edit” model editing algorithms, ROME (Meng et al., 2023a), MEMIT (Meng et al., 2023b), PMET (Li et al., 2024) and AlphaEdit (Fang et al., 2025), and applied to four LLMs: GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021), Llama-2 (7B) (Touvron et al., 2023), and Llama-3.1 (8B) (Meta, 2024). Our experiments are conducted on two widely used factual knowledge editing benchmarks: COUNTERFACT (Meng et al., 2023a) and ZsRE (Levy et al., 2017). Our results show that iterative model editing improves edit success while also reducing the approximation error introduced by the rank-one update. Furthermore, we demonstrate that incorporating even a single neighboring knowledge during model editing reduces unintended modifications to neighboring knowledge at test time, resulting in stronger edit performance. Overall, our proposed methods are broadly applicable and consistently effective across current locate-and-edit approaches that adopt the two-stage editing framework, and are readily applicable to future methods built on this foundation.
@@ 2 The Locate-and-Edit Framework
Prior work focuses on editing the factual knowledge within the LLM. Factual knowledge is represented as a triplet $(s,r,o)$, where subject $s\in\mathcal{X}^{T_{s}}$, relation $r\in\mathcal{X}^{T_{r}}$, and object $o\in\mathcal{X}$ are sequences of tokens, e.g., The iPad $[s]$ is produced by $[r]$ Apple $[o]$. We consider only single token objects in this representation, following previous work Meng et al. (2023a, b); Li et al. (2024). The model editing task is to make the model place a higher likelihood on a new object $o^{*}$ than an old object $o$ when presented with $x=(s,r)$, i.e., find new parameters $\theta^{\prime}$, such that $f_{\theta^{\prime}}(x)[o^{*}]>f_{\theta^{\prime}}(x)[o]$. Model editing is not limited to a single edit, but can encompass a batch of $m$ desired edits $D=\{(s_{i},r_{i},o_{i},o^{*}_{i})\}_{i=1}^{m}$.
Locate and edit model editing algorithms hypothesize that factual knowledge locates within specific layers of the LLM, and updating parameters in these layers is sufficient to induce the desired change in object (Pearl, 2013; Vig et al., 2020; Meng et al., 2023a) . These methods employ causal tracing to identify these layers responsible for the factual knowledge, referred to as causal layers $\{l_{1},\ldots,l_{c}\}$, where $c$ denotes the number of layers in this range. The last MLP in these layers has been found to have a major impact on the object token distribution when presented with the subject tokens (Meng et al., 2023a, b; Geva et al., 2021). Due to this major impact, locate and edit methods focus on only updating these MLP weights.
The weight update is performed in two stages: Optimization stage finds ideal values for the network’s hidden state in certain transformer layer to make $o^{*}$ likely and Spread stage updates the weights of the last MLP in casual layer(s) to approximate this ideal hidden state. We detail these stages below for MEMIT and discuss its differences to PMET, AlphaEdit, and ROME.
@@
MEMIT (Meng et al., 2023b) $h^{l_{c}}$ (hidden state) $W^{l_{1}}\cdots W^{l_{c}}$ (least-squares update)
PMET (Li et al., 2024) $\textit{Attn}^{l_{c}}$ $\textit{z}^{l_{c}}$ (ideal attention + MLP) $W^{l_{1}}\cdots W^{l_{c}}$ (attention-free update)
@@
ROME (Meng et al., 2023a) $z^{l_{*}}$ (ideal value) $W^{l_{*}}$ (rank-one equality-constrained update)
@@
The ideal hidden state for the prompt $x=(s,r)$ is $\bar{h}_{t}^{l_{c}}=h_{t}^{l_{c}}+\delta^{*}$, where $\delta^{*}$ is found by performing gradient descent on $g$. This ideal hidden state is then used in computing weight update in the next stage. AlphaEdit is the same as MEMIT, whereas PMET differs from MEMIT by searching for ideal outputs for the attention module and MLP modules in layer $l_{c}$. ROME searches for an ideal output for the MLP module of a single layer in the set of causal layers.
@@
Several other locate-and-edit methods exist, such as R-ROME (Gupta et al., 2024a), EMMET (Gupta et al., 2024c), EVOKE (LTI) (Zhang et al., 2025), and ENCORE (Gupta et al., 2025), all of which follow the same two-stage procedure (detailed in Appendix A). Below, we discuss the fundamental limitations of this framework and propose a general approach that applies broadly to all methods in this category—not just the four highlighted here, as summarized in Table 1.
@@ 3.1 Iterative Model Editing
There are two possible reasons for UnderEdit to occur. The first is that $\bar{h}^{l_{c}}_{t}$ does not reflect a hidden state for a successful edit. The second is that the weight update results in $\hat{h}^{l_{c}}_{t}\not\approx\bar{h}^{l_{c}}_{t}$. We hypothesize that both of these potential problems can be addressed by running the memory edit process multiple times because: 1) it allows for potentially finding better $\bar{h}^{l_{c}}_{t}$ after updating the model parameters so that $\|\hat{h}_{t}^{l_{c}}-\bar{h}^{l_{c}}_{t}\|\leq\|h^{l_{c}}_{t}-\bar{h}^{l_{c}}_{t}\|$, and 2) on the next iteration, the approximation used in the Spread stage for the weight update will be better since $\hat{h}^{t}_{l_{c}}$ is closer to $\bar{h}_{t}^{l_{c}}$ than $h_{t}^{l_{c}}$. We detail this iterative process below for MEMIT, but it can also be adapted to ROME, PMET and AlphaEdit by replacing $\bar{h}_{t}^{l_{c}}$ with the targets of the optimization procedure for those algorithms.
@@
For brevity, we use $\Delta p_{k}$ to denote the above difference in perplexity in iteration $k$. Empirically, we found $\epsilon=1$ to be a sufficient threshold for the data sets used in this paper.
@@ 3.2 Neighbor-Assisted Model Editing
Model editing must not only change the model’s output from $o$ to $o^{*}$ given $(s,r)$, but also preserve outputs for neighboring knowledge, i.e $(\tilde{s},r,o)$, where $\tilde{s}$ is a new subject sharing the same relation $r$. A preservation example is shown in Figure 1a: iPhone 11 [$\tilde{s}$] is still produced by [$r$] Apple [$o$] despite iPad is edited to produced by Honda [^1] In the COUNTERFACT dataset.
@@
Table 2: Iterative model editing results on COUNTERFACT for at most 10 iterations (denoted by k). We compare the evaluation metrics of iteration that met stopping criterion $|\Delta p_{k}|\leq 1$ to that of their corresponding first iteration and bold the higher value. PMET on GPT-2 XL require more than 5 iterations to achieve our stopping criteria. Results for all iterations are provided in Table 9 (Appendix G). While ROME is known to collapse (results reported in Table 9), we observed a unique case of collapse with Llama-2 (7B) specifically when using MEMIT. We discuss this in Section 5. Note: Results for PMET on LLaMA-3.1 are based on a single subset, as results for the remaining subsets are still being generated. Results for AlphaEdit on GPT-J are currently unavailable due to an implementation error in EasyEdit, which we are actively working to resolve.
@@ 4.1 Datasets
To evaluate model editing across different datasets, we use the COUNTERFACT Meng et al. (2023a) and ZsRE Levy et al. (2017) datasets. Both datasets consist of approximately 20k factual knowledge instances. Due to hardware limitations, for each model-editing experiment, we ran it on a subset of $m=1,\!000$ edits for each dataset. We repeat the editing task three times each using a different set of $m$ edits sampled from the whole dataset. We ensured that the edits in these trials were mutually exclusive and report the averages across them.
@@
4.2 Evaluation Metrics
We are primarily concerned with evaluating how well iterative and neighbor-assisted model editing reduce the frequency of UnderEdit and OverEdit. We measure how successful the editing algorithms were by examining efficacy and generalization scores. Efficacy measures the success of introducing new knowledge edits in the dataset. Generalization tests whether the edit is robust and not overfit by evaluating the model on paraphrases of the examples in the dataset. To understand how well the algorithms were at avoiding OverEdits, we measure the specificity of the model, i.e., how much of the neighbor knowledge remained unchanged. To summarize the overall performance in a score, we use a harmonic mean of efficacy, generalization, and specificity.
@@ 5.1 Iterative Model Editing Results
We conducted iterative model editing experiments across all datasets, LLMs, and editing algorithms, running each configuration for at most 10 iterations. The evaluation results are presented in Table 2 for COUNTERFACT and Table 10 for ZsRE (provided in Appendix 10 due to space constraints). From these experiments results, we drew several conclusions.
@@
Although iterative editing is effective in most cases, we also observed model collapse as editing progresses, indicated by high model perplexity on ME-PPL-50 (5 out of 15 experiment settings). This collapse behavior aligns with the continuous editing failures observed in previous work Gupta et al. (2024b); Meng et al. (2023a). This suggests that when the combination of the model and editing algorithm succeeds in continuous editing, an essential experimental setting, iterative editing can further improve model performance. Specifically, ROME collapses under iterative 333 ROME does not support batch editing, as it can only modify one fact at a time Meng et al. (2023b). We discuss this further in Appendix A. editing, which aligns with prior findings on its inability to support continuous editing Gupta et al. (2024b). We also found that Llama-2 (7B) collapses only when edited with MEMIT, a result consistent with findings in Yang et al. (2024). Thus, we conclude that iterative model editing does not inherently lead to collapse. However, unstable models that fail with sequential editing may not benefit from this approach. Although investigating collapse is not the primary focus of this work, our findings (detailed in Appendix E) offer a useful foundation for future research in this area.
We also consistently observed a trade-off between generalization and specificity in Table 2. To further investigate this phenomenon, we focused on AlphaEdit, which aims to address this trade-off using null space projection to constrain model updates so that they minimally interfere with existing knowledge. Our results show that although AlphaEdit is designed to mitigate OverEdit, its performance still degrades in specificity under iterative model editing—indicating that this method does not fully resolve the problem. This persistent trade-off affects all locate-and-edit algorithms, and the proposed iterative editing is effective and widely applicable across this family of methods.

### v3

@@ Abstract
Large Language Models (LLMs) are widely deployed in downstream tasks, but keeping their knowledge up-to-date via retraining or fine-tuning is often computationally expensive. Model editing provides a more efficient alternative by updating a targeted subset of parameters, which often follows the locate-and-edit paradigm. Despite this efficiency, existing methods are limited: edits may fail to inject knowledge (UnderEdit) or unintentionally disrupt unrelated neighboring knowledge (OverEdit). To address these challenges, we propose two complementary methods: iterative model editing, which applies successive edits to mitigate UnderEdit, and neighbor-assisted model editing, which incorporates neighboring knowledge during editing to reduce OverEdit. Our extensive experiments show that these techniques improve editing performance across multiple LLMs, algorithms, and benchmarks, reducing UnderEdit by up to 38 percentage points and OverEdit by up to 6, while remaining broadly applicable to any locate-and-edit method. We release our code at https://github.com/bhimanbaghel/ResolveUnderOverEdit.
@@
Iterative & Neighbor-Assisted Model Editing
Bhiman Kumar Baghel   Emma Jordan   Zheyuan Ryan Shi   Xiang Lorraine Li Department of Computer Science, University of Pittsburgh, PA, USA {bkb45, xianglli}@pitt.edu
@@ 1 Introduction
LLMs have been widely used as repositories of factual and specialized knowledge (Petroni et al., 2020; Jiang et al., 2021; Roberts et al., 2020; Youssef et al., 2023). However, the world is constantly changing, with knowledge and information evolving rapidly, such as significant government policy changes and their wide impacts across various domains. Thus, it is essential for many NLP applications, such as text generation, question answering, and knowledge retrieval, to have models that can adapt to knowledge changes both effectively and efficiently. Re-training an LLM is resource-intensive (Patterson et al., 2021). Standard supervised fine-tuning is data hungry and less effective (Meng et al., 2023). Model-editing, which directly modifies important model parameters for making the prediction, has emerged as a more efficient alternative for updating outdated information (Meng et al., 2022, 2023; Li et al., 2024; Fang et al., 2025). These methods adopt a “locate-and-edit” approach, where they first identify the parameter locations associated with outdated knowledge and then update the parameters to enable the model to incorporate and predict the new knowledge.
@@
In summary, we propose solutions to two fundamental challenges in model editing: UnderEdit, where edits fail, and OverEdit, where neighboring knowledge is erroneously modified. We evaluate our approach using four “locate and edit” model editing algorithms, ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), PMET (Li et al., 2024) and AlphaEdit (Fang et al., 2025), and applied to four LLMs: GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021), Llama-2 (7B) (Touvron et al., 2023), and Llama-3.1 (8B) (Meta, 2024). Our experiments are conducted on two widely used factual knowledge editing benchmarks: COUNTERFACT (Meng et al., 2022) and ZsRE (Levy et al., 2017). Our results show that iterative model editing improves edit success while also reducing the approximation error introduced by the rank-one update. Furthermore, we demonstrate that incorporating even a single neighboring knowledge during model editing reduces unintended modifications to neighboring knowledge at test time, resulting in stronger edit performance. Overall, our proposed methods are broadly applicable and consistently effective across current locate-and-edit approaches that adopt the two-stage editing framework, and are readily applicable to future methods built on this foundation.
@@ 2 The Locate-and-Edit Framework
Prior work focuses on editing the factual knowledge within the LLM. Factual knowledge is represented as a triplet $(s,r,o)$, where subject $s\in\mathcal{X}^{T_{s}}$, relation $r\in\mathcal{X}^{T_{r}}$, and object $o\in\mathcal{X}$ are sequences of tokens, e.g., The iPad $[s]$ is produced by $[r]$ Apple $[o]$. We consider only single token objects in this representation, following previous work Meng et al. (2022, 2023); Li et al. (2024). The model editing task is to make the model place a higher likelihood on a new object $o^{*}$ than an old object $o$ when presented with $x=(s,r)$, i.e., find new parameters $\theta^{\prime}$, such that $f_{\theta^{\prime}}(x)[o^{*}]>f_{\theta^{\prime}}(x)[o]$. Model editing is not limited to a single edit, but can encompass a batch of $m$ desired edits $D=\{(s_{i},r_{i},o_{i},o^{*}_{i})\}_{i=1}^{m}$.
Locate and edit model editing algorithms hypothesize that factual knowledge locates within specific layers of the LLM, and updating parameters in these layers is sufficient to induce the desired change in object (Pearl, 2013; Vig et al., 2020; Meng et al., 2022). These methods employ causal tracing to identify these layers responsible for the factual knowledge, referred to as causal layers $\{l_{1},\ldots,l_{c}\}$, where $c$ denotes the number of layers in this range. The last MLP in these layers has been found to have a major impact on the object token distribution when presented with the subject tokens (Meng et al., 2022, 2023; Geva et al., 2021). Due to this major impact, locate and edit methods focus on only updating these MLP weights.
The weight update is performed in two stages: Optimization stage finds ideal values for the network’s hidden state in certain transformer layer to make $o^{*}$ likely and Spread stage updates the weights of the last MLP in casual layer(s) to approximate this ideal hidden state. We detail these stages below for MEMIT (Meng et al., 2023) and discuss its differences to PMET (Li et al., 2024), AlphaEdit (Fang et al., 2025) and ROME (Meng et al., 2022).
@@
MEMIT (Meng et al., 2023) $h^{l_{c}}$ (hidden state) $W^{l_{1}}\cdots W^{l_{c}}$ (least-squares update)
PMET (Li et al., 2024) $\textit{Attn}^{l_{c}}$ & $\textit{z}^{l_{c}}$ (ideal attention + MLP) $W^{l_{1}}\cdots W^{l_{c}}$ (attention-free update)
@@
ROME (Meng et al., 2022) $z^{l_{*}}$ (ideal value) $W^{l_{*}}$ (rank-one equality-constrained update)
@@
The ideal hidden state for the prompt $x=(s,r)$ is $\bar{h}_{t}^{l_{c}}=h_{t}^{l_{c}}+\delta^{*}$, where $\delta^{*}$ is found by performing gradient descent on $g$. This ideal hidden state is then used in computing weight update in the next stage. AlphaEdit is the same as MEMIT in this stage, whereas PMET differs from MEMIT by searching for ideal outputs for the attention module and MLP modules in layer $l_{c}$. ROME searches for an ideal output for the MLP module of a single layer in the set of causal layers.
@@
Building on the preceding discussion of the primary methods, several additional locate-and-edit algorithms are instances of the same two-stage framework: R-ROME (Gupta et al., 2024a), which conditions on more context prompts than ROME to improve update stability; EMMET (Gupta et al., 2024c), which extends ROME to batch edits; EVOKE (LTI)(Zhang et al., 2025), which adds constraints in the Optimization stage to regularize the latent search; and ENCORE(Gupta et al., 2025), which uses early stopping in Optimization to limit overconfident hidden states and normalized updates in Spread to control update magnitude, thereby mitigating overconfident editing and instability. Table 1 provides an overview of how these methods implement the two stages, and Appendix A details the algorithms. We next examine inherent limitations of this two-stage framework and propose a general, method-agnostic remedy that applies broadly across such approaches.
@@ 3.1 Iterative Model Editing
There are two possible reasons for UnderEdit to occur. The first is that $\bar{h}^{l_{c}}_{t}$ does not reflect a hidden state for a successful edit. The second is that the weight update results in $\hat{h}^{l_{c}}_{t}\not\approx\bar{h}^{l_{c}}_{t}$. We hypothesize that both of these potential problems can be addressed by running the memory edit process multiple times because: 1) it allows for potentially finding better $\bar{h}^{l_{c}}_{t}$ after updating the model parameters so that $\|\hat{h}_{t}^{l_{c}}-\bar{h}^{l_{c}}_{t}\|\leq\|h^{l_{c}}_{t}-\bar{h}^{l_{c}}_{t}\|$, and 2) on the next iteration, the approximation used in the Spread stage for the weight update will be better since $\hat{h}^{t}_{l_{c}}$ is closer to $\bar{h}_{t}^{l_{c}}$ than $h_{t}^{l_{c}}$. Our approach involves whole-pipeline iteration—executing Optimization followed by Spread across multiple iterations until convergence. In contrast, editing algorithms (Table 1) perform optimization only once within the Optimization stage using gradient descent, which finds the target representation but does not by itself yield a better edit. Our proposed iterative refinement ensures that both the target representation and the subsequent parameter update improve in tandem. We detail this iterative process below for MEMIT, but it can also be adapted to ROME, PMET and AlphaEdit by replacing $\bar{h}_{t}^{l_{c}}$ with the targets of the optimization procedure for those algorithms.
@@
For brevity, we use $\Delta p_{k}$ to denote the above difference in perplexity in iteration $k$. Empirically, we found $\epsilon=1$ to be a sufficient threshold for the data sets used in this paper. We set $\epsilon=1$ because, across all datasets considered, this threshold consistently delivered the best (or statistically indistinguishable) performance while minimizing the number of iterations (Table 2). Table 9 further justifies this choice, showing that smaller thresholds yield negligible gains at substantially higher compute cost, whereas larger thresholds can degrade accuracy.
The algorithm aims to reduce the perplexity difference between the two stages mainly because we found a mismatch between these two perplexity values. The Optimization stage seeks to learn a hidden representation $\bar{h}_{t,k}^{l_{c}}$ that minimizes the perplexity of the target token, however, we found that after Spread this perplexity increases, leading to UnderEdit (as shown in Table 9 and 11). Thus we directly minimize this discrepancy and thereby mitigate UnderEdit. We provide further empirical verification of this relation in Appendix F.1.
@@ 3.2 Neighbor-Assisted Model Editing
Model editing must not only change the model’s output from $o$ to $o^{*}$ given $(s,r)$, but also preserve outputs for neighboring knowledge, i.e $(\tilde{s},r,o)$, where $\tilde{s}$ is a new subject sharing the same relation $r$. A preservation example is shown in Figure 1a: iPhone 11 $\tilde{s}$] is still produced by [$r$] Apple [$o$] despite iPad is edited to produced by Honda [^1] In the COUNTERFACT dataset.
@@
Table 2: Iterative model editing results on COUNTERFACT for at most 10 iterations (denoted by k). We compare the evaluation metrics of iteration that met stopping criterion $|\Delta p_{k}|\leq 1$ to that of their corresponding first iteration and bold the higher value. PMET on GPT-2 XL require more than 5 iterations to achieve our stopping criteria. Results for all iterations are provided in Table 9 (Appendix G). While ROME is known to collapse (results reported in Table 9), we observed a unique case of collapse with Llama-2 (7B) specifically when using MEMIT. We discuss this in Section 5. Note: Unlike EasyEdit Wang et al. (2024b), which reports a single-edit baseline, batch editing with ROME is performed to ensure consistency with the setups of other algorithms.
@@ 4.1 Datasets
To evaluate model editing across different datasets, we use the COUNTERFACT Meng et al. (2022) and ZsRE Levy et al. (2017) datasets. Both datasets consist of approximately 20k factual knowledge instances. Due to hardware limitations, for each model-editing experiment, we ran it on a subset of m=1,000m=1,\!000 edits for each dataset. We repeat the editing task three times each using a different set of $m$ edits sampled from the whole dataset. We ensured that the edits in these trials were mutually exclusive and report the averages across them.
@@
4.2 Evaluation Metrics
We are primarily concerned with evaluating how well iterative and neighbor-assisted model editing reduce the frequency of UnderEdit and OverEdit. We measure how successful the editing algorithms were by examining efficacy and generalization scores. Efficacy measures the success of introducing new knowledge edits in the dataset. Generalization tests whether the edit is generalizable by evaluating the model on paraphrases of the examples in the dataset. To understand how well the algorithms were at avoiding OverEdits, we measure the specificity of the model, i.e., how much of the neighbor knowledge remained unchanged. To summarize the overall performance in a score, we use a harmonic mean of efficacy, generalization, and specificity.
@@ 5.1 Iterative Model Editing Results
We conducted iterative model editing experiments across all datasets, LLMs, and editing algorithms, running each configuration for at most 10 iterations. The evaluation results are presented in Table 2 for COUNTERFACT and Table 10 for ZsRE (provided in Appendix due to space constraints). From these experiments results, we drew several conclusions.
@@
Although iterative editing is effective in most cases, we also observed model collapse as editing progresses, indicated by high model perplexity on ME-PPL-50 (6 out of 16 experiment settings). This collapse behavior aligns with the continuous editing failures observed in previous work Gupta et al. (2024b); Meng et al. (2022). This suggests that when the combination of the model and editing algorithm succeeds in continuous editing, an essential experimental setting, iterative editing can further improve model performance. Specifically, ROME collapses under iterative  [^3] ROME does not support batch editing, as it can only modify one fact at a time Meng et al. (2023). We discuss this further in Appendix A. editing, which aligns with prior findings on its inability to support continuous editing Gupta et al. (2024b). We also found that Llama-2 (7B) collapses only when edited with MEMIT, a result consistent with findings in Yang et al. (2024). Thus, we conclude that iterative model editing does not inherently lead to collapse. We further establish this is Appendix E. However, unstable models that fail with sequential editing may not benefit from this approach. Although investigating collapse is not the primary focus of this work, our findings offer a useful foundation for future research in this area.
We also consistently observed a trade-off between generalization and specificity in Table 2. Specifically, we focused on AlphaEdit, which aims to address this trade-off using null space projection to constrain model updates so that they minimally interfere with existing knowledge. Our results show that although AlphaEdit is designed to mitigate OverEdit, its performance still degrades in specificity under iterative model editing—indicating that this method does not fully resolve the problem. This reflects a No-Free-Lunch effect within locate-and-edit methods—improving one objective (e.g., efficacy) inevitably harms another (e.g., specificity). This persistent trade-off affects all locate-and-edit algorithms, and the proposed iterative editing is effective and widely applicable across this family of methods.